OpenStudy (anonymous):

how to solve 15x-10=-9x+2

OpenStudy (hba):

Well Its Very Easy To Solve,Why Dont you give it a try ..

OpenStudy (nubeer):

first bring all the terms with "x" on one side and terms without it on the other side.. Remember when u move any number from one side of equal to other side u have to change its sign for example: +2x+ 1 = 3x now u have to bring all terms with x on one side. 1 = 3x-2x so can u give it a try?

OpenStudy (anonymous):

my answer is x=1

Parth (parthkohli):

I'm going to make up an equation and show all steps. \[ \begin{aligned}3x + 2& = 9x + 3 \\ 3x&=9x + 3 - 2\\3x&=9x + 1\\ 3x-9x&=1 \\ -6x&=1\\x &={-1 \over 6}\end{aligned}\]

OpenStudy (nubeer):

well it's not 1.. can u do all the steps here.. so we can see where u made the mistake.

OpenStudy (anonymous):

thxs i will re work

OpenStudy (anonymous):

15x-10=9x+2 15x=9x+2-10 5x=9x+-8 5x-9x=-8 -4x=-8 x=-2

OpenStudy (anonymous):

pls chk my ans

OpenStudy (nubeer):

somehow in 3rd line u made 15x into 5x .. lol u need to see it again. 15x-10=9x+2 this is not same as the question u posted above..